OpenStudy (anonymous):

4 times the measure of the complement of an angle is to equal to 2/3 the measure of the original angle's supplement. what is the measure of the angle?

OpenStudy (anonymous):

i neeeeed help!

OpenStudy (raden):

let the original angle is x. according information above the expression will be 4(90-x) = (2/3)x solve for x

OpenStudy (charlotte123):

anyone help me with my question? :) its easy and will get a medal for trying ^-^

OpenStudy (raden):

wooppss misread, that's original supplement :) repeat again : 4(90-x) = (2/3)(180-x)

OpenStudy (anonymous):

im so lost now..

OpenStudy (raden):

well, look how the algebraic principle works 4(90-x) = (2/3)(180-x) times 3 from both sides, first